# Realistic Cost to Execute Practical Quantum Circuits using Direct Clifford+T Lattice Surgery Compilation

### v2

@@ 2.2. Overview of Compilation Strategy
As mentioned in Sec. 1, we choose Clifford+T for our input logical instruction set and do not perform any global circuit re-write. Within the scope of this paper, it is assumed that the compilation of a quantum algorithm into a logical circuit has already taken place. For an output surface code instruction set, we have chosen a local lattice surgery instruction set based upon the one from Ref. (Beverland et al., 2022a) but with refinements from Ref. (LeBlond et al., 2023a). We also make use of an intermediate representation that allows us to translate logical gates into an abstract instruction set within the fault-tolerant layer before routing and ordering operations into logical time-slices (slicing) takes place. In this section we will provide an overview of all three of these instruction sets as well as our choice of layout.
@@ 2.2.1. Circuit-Level Logical Instruction Set
Here, we will briefly describe the input gate set that we have in mind and overview the approach to compilation for each gate. We begin with a logical circuit expressed using the Clifford+T gate set (more specifically, we allow X, Y, Z, S, $\mathrm{S}^{\dagger}$, T, $\mathrm{T}^{\dagger}$, CNOT, H, and Reset gates). Pauli gates (X, Y, and Z) are not compiled into any explicit operations because they can be tracked and later combined with measurements according to the Pauli frame tracking methodology (Fowler et al., 2012; Riesebos et al., 2017) [^9] We do not include functionality for tracking the Pauli frame from this information because it has no bearing on resource estimation, which is our ultimate goal for compilation. Besides, the actual run-time Pauli frame will also depend on measurement outcomes gathered during circuit execution.. We compile both S/$\mathrm{S}^{\dagger}$ and T/$\mathrm{T}^{\dagger}$ gates using standard gate teleportation circuits from Ref. (Fowler et al., 2012). The S gate teleportation circuit that we use is “catalytic” in the sense that Y states input to the protocol are re-produced by the circuit [^10] In the ensuing analysis we will neglect the resource cost of Y state generation because we assume it to be negligible.. Because S and $\mathrm{S}^{\dagger}$ gates utilize the same circuit and the only difference relies on a Pauli Z correction (which is not explicitly applied), we treat them the same within our compiler. Similarly, we treat T/$\mathrm{T}^{\dagger}$ the same because they utilize the same teleportation circuit except for the application of a corrective S gate, which we always explicitly compile. In opposition to the streaming philosophy underlying the original LSC (Watkins et al., 2023), we conceive of the whole circuit being compiled in advance of execution in order to aid in our efforts toward parallelism. Thus, because in the cases of both T and $\mathrm{T}^{\dagger}$ the corrective gate is required 50% of the time, we always allocate the required active volume in our compilation and imagine that the system executing the compiled program uses run-time measurement outcomes to apply corrective operations or not [^11] While this may seem like a downside, this choice significantly improves our ability to detect and compile parallel operations (see Sec. 2.3) and optimize resources dedicated to magic state distillation and storage (see Sec. 4).. As mentioned, CNOT gates, which these gate teleportation also rely on, are performed using the Bell-based protocol of Ref. (Beverland et al., 2022a) (to be discussed). Lastly, H gates are always followed by a patch rotation in order to preserve the orientation of patches’ logical X and Z operators throughout the computation, which allows us to comply with the definition of logical tile offered in Ref. (LeBlond et al., 2023a).
@@ 2.2.2. Gate-by-gate Processing into Intermediate Representation
The first stage of lattice surgery compilation maps the Clifford+T gate set into an abstract instruction set which acts as an intermediate representation within the fault-tolerant layer. This instruction set is not to be confused with the local lattice surgery instruction set that will compose the final compiled output; the key difference between them is that the intermediate representation acts on logical qubit (patch) IDs while the local instructions act on tiles. Being an extension of what were called logical lattice instructions (LLI) in Ref. (Watkins et al., 2023), we will continue use the term LLI to refer to this intermediate representation for the sake of continuity. Due to the nature of lattice surgery compilation requiring the preparation of ancillae and usage of resource states that were not present in the original logical circuit, new patch IDs need to be generated throughout this stage of compilation. These patch IDs remain virtual until bound with patches instantiated on tiles in the second stage of compilation. It is worth noting here that only a logical dependency graph of LLI results from the first stage of compilation, while the product of the second stage is LLI organized onto logical time-slices (sliced LLI) with their compilation into local instructions given in brackets (if applicable). The former depends on dependencies of LLI alone while the latter depends upon the availability of routes, resource states, and the busy status of patches.
@@
We will now overview how each of the logical gates mentioned in the previous section map into LLI (or sequences of LLI) in the first stage of compilation. Pauli X, Y, and Z are themselves treated as LLI. CNOT gates are simply compiled into BellBasedCNOT LLI, which act on the control and target patch IDs. We note that, while local compilation of the BellBasedCNOT LLI requires the preparation and subsequent measurement of many new logical qubit patches, we do not assign patch IDs to them because they do not have any impact on the logical dependency graph and thus their assignment is unnecessary. H gates are themselves treated as LLI but are followed by a RotateSingleCellPatch LLI. S/$\mathrm{S}^{\dagger}$ gates begin with a YStateRequest LLI acting on a newly generated patch ID (later to be associated with the nearest patch containing a Y state). Then, the appropriate sequence of LLI to execute the required H and CNOT gates according to the catalytic S gate circuit in Ref. (Fowler et al., 2012) are added. Lastly, T/$\mathrm{T}^{\dagger}$ gates begin with a MagicStateRequest LLI acting on a newly generated patch ID (later to be associated with the nearest patch containing a magic state) followed by the BellBasedCNOT LLI and SinglePatchMeasurement LLI required to implement the gate teleportation circuit from Ref. (Fowler et al., 2012). The LLI required by the corrective S gate then follow.
@@ 2.2.3. Further Processing of Intermediate Representation
The second stage of lattice surgery compilation involves the routing and slicing of the LLI resulting from the first stage. In order to translate the position-agnostic intermediate representation into instructions acting on hardware tiles, the compiler must allocate an appropriate space-time volume to each instruction while avoiding resource conflicts. The layout, which is a configurable input to the compiler (Fig. 1), provides a set of constraints at the outset of compilation. Subsequently processed LLI modify these constraints by instantiating patches on tiles along routes, binding or un-binding resource states with IDs, and/or changing the active status of patches and their edges during lattice surgery operations. Once no further LLI can be compiled onto a slice the compiler advances to the next slice, which can also modify the constraints by e.g. re-setting tiles after measurements or re-supplying magic states according to the compilation settings. Given our input instruction set and intermediate representation, our compiler binds patch IDs to available magic states and Y states, allocates busy tiles for RotateSingleCellPatch LLI, and compiles BellBasedCNOT LLI into local lattice surgery instructions in the second compilation stage.
Figure 2. The path corresponding to a CNOT.
Figure 2. The path corresponding to a CNOT.
Figure 3. Two CNOTs avoiding each other.
We use results from Ref. (Beverland et al., 2022a) to compile BellBasedCNOT LLI into local instructions. In that protocol, the preparation and usage of a long-range Bell state requires the usage of every tile located along a free route between the control and target patches for two logical time-slices (see Fig. 3 for an example of a BellBasedCNOT route), which involves two layers of lattice surgery operations to produce local Bell states and perform Bell measurements between their end-points [^12] The specific local instructions that the BellBasedCNOT are compiled to depend on whether there is an even or odd number of tiles along the route between the control and target patches, according to Fig. 19a and c in Ref. (Beverland et al., 2022a).. While this method involves more local lattice surgery operations than do CNOT compilation protocols involving extended ancilla patches (Horsman et al., 2012; Litinski and von Oppen, 2018), it does not have increased resource overhead relative to them. Implementation details for this protocol will be given in Sec. 2.3. We re-iterate that our intention in compiling to a local instruction set rather than using extended ancilla patches is that it facilitates a simple reduction into primitives that allows for easier end-to-end compilation and circuit verification (LeBlond et al., 2023a).
@@ 2.2.4. Contention for Space
Operations that are logically parallel may not be parallelizable within the fault-tolerant layer due to geometrical constraints. For example, a CNOT gate between two qubits requires an ancilla path between the control and target, a T gate requires an ancilla path from a logical qubit to a magic state, and a Hadamard gate requires an uninitialized neighboring ancilla tile so that a patch rotation may be performed. See Fig. 3 for a simple example of resource contention between two CNOT gates. Ref. (Beverland et al., 2022a) has proposed one method for retaining logical parallelism in the fault-tolerant layer by finding maximal edge disjoint path sets. Among other things, they found that a layer of $k$ parallel CNOT gates needs depth $\Omega(\sqrt{k})$ to execute on a surface code architecture in the worst-case. In our compilation strategy, we exploit logical parallelism within the fault-tolerant layer by relying on a simple heuristic algorithm (A*) to find vertex disjoint paths and greedily allocating these paths to slices (see Algorithm 1 in Sec. 2.3 for details). In Sec. 3 we will show that, in our implementation, the number of slices scales similarly to the expectations from Ref. (Beverland et al., 2022a).
@@ 2.2.5. Choice of Surface Code Layout
Within a lattice surgery paradigm, surface code patches can be simplistically thought of as rectangular entities that are merged and split through edge-wise interactions to generate entanglement or transport quantum information. This results in a type of quantum circuit layout problem, where in general the layout of qubits on a constrained hardware architecture is optimized to reduce circuit depth (Paler et al., 2023). In surface code compilation, layouts of equally-sized tiles on a two-dimensional grid are often assumed (Litinski, 2019a; Beverland et al., 2022a), though it is not always beneficial to have only have one tile size available throughout a computation (Litinski, 2019b; Beverland et al., 2022b). Equal X and Z code distances are also not necessarily beneficial in the realistic setting of biased noise (Chamberland and Campbell, 2022b). In proposals for quantum computer architectures, there is a trade-off between compression of layouts in space and the accessibility of surface code patch edges, as compression in space can lead to overhead in time due to the need to rotate patches to expose needed operators (Litinski, 2019a). While keeping the number of tiles dedicated to routing space minimal under the condition that all patch edges are exposed is generally considered to be desirable in sequential Pauli-based computation (Litinski, 2019a; Beverland et al., 2022b; Chamberland and Campbell, 2022b), it is not as clear in the case of direct Clifford+T compilation where liberal routing space can aid in parallel compilation of logically parallel gates (Beverland et al., 2022a). In this paper, we take the definition of logical tile presented in Ref. (LeBlond et al., 2023a) and make use of a family of layouts related to the one from Ref. (Beverland et al., 2022a) [^13] We will call this family of layouts EDPC layouts because of the paper that proposed them.. We assume that tile sizes are uniform within lattice surgery compilation since, while we do use two levels of distillation with two different code distances, we do not explicitly compile distillation circuits (as explained in Sec. 4).
@@ 2.3. Our Revisions to the Lattice Surgery Compiler
Much of the functionality needed to implement our compilation strategy from Sec. 2.2 did not originally exist in LSC. However, we leveraged the extensibility of the LSC framework to improve layout generation including the handling of magic states, expand the options for processing logical gates into the intermediate representation (LLI), create a compilation pipeline to handle DAG-based parallelism in the LLI layer, and implement a new compilation layer for processing LLI into local instructions. In this section, we will provide details about our revisions to the LSC in a few areas where we feel that more specificity is warranted than what was presented in Sec. 2.2. We will first review some of the functionality of the original LSC from Ref. (Watkins et al., 2023) in order to provide context for our contributions to it.
@@ 2.3.1. Lattice Surgery Compiler Review
  • Compilation is implemented as a streamed process, so that compilation could be performed in real-time based on input to classical control software while the quantum computer executes previously compiled instructions. The streaming pipeline (which we retained as an option in our revised version) has the benefit of being able to compile circuits that are too large to fit in memory. It also allows for a limited form of parallelism by compiling sequential instructions onto the same slice until it encounters one that cannot be routed. However, the parallelism is limited by the lack of a logical dependency graph prior to time-slicing instructions, and by multi-stage instructions which automatically request new time-slices for follow-up instructions.
@@ 2.3.2. Layout Generation
We have improved the layout generation capabilities within the LSC by implementing three new types of tiles (specified by text characters as described previously). In the new version of the compiler, there are specified tiles that are reserved for magic states (labeled ‘M’) on which magic states are replenished according to the disttime and nostagger parameters, which respectively specify the rate at which magic states are replenished at these tiles and whether or not they are replenished in a staggered fashion. This improvement enables compilation of T/$\mathrm{T}^{\dagger}$ gates to be dissociated from the space-time footprint of magic state distillation, allowing much greater flexibility in resource estimation and optimization. The post-hoc resource analysis discussed in Sec. 4 accounts for both the space-time footprint of magic state factories and the cost to maintain distilled state storage banks, both of which are determined after the cadence of sliced MagicStateRequest LLI is known from the LSC output. Although magic state distillation is not specifically compiled in our framework, our analysis of magic state management (generation, storage, and consumption) makes no approximations as to the availability and consumption of these states at any point in the computation. The second new type of tile we have implemented are tiles that store pre-distilled Y states (labeled ‘Y’). (Previously, they were produced on-the-fly according to a twist-based protocol (Watkins et al., 2023)). Finally, dead tiles (tiles on which no operations can be performed and which are avoided by the router) have also been implemented (labeled ‘X’) so that non-rectangular layouts can be specified (though we have not yet made use of this).
@@
For our resource estimates, we implemented a layout generator that constructs the EDPC layout as illustrated in Figure 4, accessible through the -L edpc command line option (Fig. 4(a) represents the EDPC layout that is generated by default using this setting). In Figure 4, the brown tiles in the bulk (bound to patch IDs as shown) correspond to patches encoding logical qubits from the input circuit. Brown and pink tiles on the boundary correspond to pre-distilled Y states and tiles that are reserved for magic states, respectively. While these have no patch IDs by default, they are bound with IDs upon encountering YStateRequest and MagicStateRequest LLI during the second stage of compilation. White tiles indicate ancilla tiles that can be used for routing (in the case of our compilation strategy, this is where long-range Bell states are prepared and where patch rotations take place). In order to observe the effects of different data:ancilla tile ratios, we implemented num_lanes and condensed options to -L edpc. The effects of these options on the generated layout are depicted in Fig. 4(b) and (c), respectively. The former option can be used to open up more lanes for routing CNOT gates at the expense of greater spatial overhead, while the latter option decreases the total routing space by blocking off two edges of each logical qubit. The effects of these options on circuit depth and space-time volume will be explored in Sec. 3.
@@
Figure 4. EDPC layouts for 9 logical qubits, generated using the (a) default, (b) num_lanes = 2, and (c) condensed settings. Numbered brown tiles in the bulk correspond to logical qubits. Brown and Pink tiles on the boundary correspond to Y states and tiles reserved for magic states, respectively. White tiles indicate tiles dedicated to routing instructions and the preparation of ancillae.

### v4

@@ 2.2. Overview of Compilation Strategy
As mentioned in Sec. 1, we choose Clifford+T for our input logical instruction set and do not perform any global circuit re-write. Within the scope of this paper, it is assumed that the compilation of a quantum algorithm into a logical circuit has already taken place. For an output surface code instruction set, we have chosen a local lattice surgery instruction set based upon the one from Ref. (Beverland et al., 2022a) but with refinements from Ref. (LeBlond et al., 2023a). We also make use of an intermediate representation that allows us to translate logical gates into a layout-independent instruction set within the fault-tolerant layer before routing and ordering operations into logical time-slices (slicing) takes place. In this section we will provide an overview of all three of these instruction sets as well as our choice of layout. See Table 1 for a summarization of our logical gate set and the associated elements of our intermediate instruction set.
@@ 2.2.1. Circuit-Level Logical Instruction Set
Here, we will briefly describe the input gate set that we have in mind and overview the approach to compilation for each gate. We allow the Clifford+T gate set; specifically, X, Y, Z, S, $\mathrm{S}^{\dagger}$, T, $\mathrm{T}^{\dagger}$, CNOT, H, and Reset gates can be used as input. The Reset gate nominally clears a logical qubit, and is simply passed directly to output without consuming any space-time volume. Pauli gates (X, Y, and Z) are treated similarly because they can be tracked in software and later combined with measurements according to the Pauli frame tracking methodology (Fowler et al., 2012; Riesebos et al., 2017) [^9] We do not include functionality for tracking the Pauli frame from this information because it has no bearing on resource estimation, which is our ultimate goal for compilation. Besides, the actual run-time Pauli frame will also depend on measurement outcomes gathered during circuit execution.. We compile both S/$\mathrm{S}^{\dagger}$ and T/$\mathrm{T}^{\dagger}$ gates using standard gate teleportation circuits from Ref. (Fowler et al., 2012). The S gate teleportation circuit that we use is “catalytic” in the sense that Y states input to the protocol are re-produced by the circuit [^10] In the ensuing analysis we will neglect the resource cost of Y state generation because we assume it to be negligible.. Because S and $\mathrm{S}^{\dagger}$ gates utilize the same circuit and the only difference relies on a Pauli Z correction (which is not explicitly applied), we treat them the same within our compiler. Similarly, we treat T/$\mathrm{T}^{\dagger}$ the same because they utilize the same teleportation circuit except for the application of a corrective S gate, which we always explicitly compile. In opposition to the streaming philosophy underlying the original LSC (Watkins et al., 2024), we conceive of the whole circuit being compiled in advance of execution in order to aid in our efforts toward parallelism and resource optimization. Thus, because in the cases of both T and $\mathrm{T}^{\dagger}$ the corrective gate is required 50% of the time, we always allocate the required active volume in our compilation and imagine that the system executing the compiled program uses run-time measurement outcomes to apply corrective operations or not [^11] While this may seem like a downside, this choice significantly improves our ability to detect and compile parallel operations (see Sec. 2.3) and optimize resources dedicated to magic state distillation and storage (see Sec. 4).. As mentioned, CNOT gates, which these gate teleportation circuits also rely on, are performed using the Bell-based protocol of Ref. (Beverland et al., 2022a) (to be discussed). Lastly, H gates are always followed by a patch rotation in order to preserve the orientation of patches’ logical X and Z operators throughout the computation, which allows us to comply with the definition of logical tile offered in Ref. (LeBlond et al., 2023a).
@@ 2.2.2. Gate-by-gate Processing into Intermediate Representation
Table 1. Circuit-level logical instruction set used in this work together with the logical lattice instructions (LLI) involved in the compilation of each instruction. Where possible, we show the numbers of tiles and slices that each gate is ultimately compiled to. Wherever BellBasedCNOTs are involved, the number of tiles is route-dependent. The details given for T/$\mathrm{T}^{\dagger}$ gates do not include those of the required corrective S gate.
Logical Gate Involved Logical Lattice Instructions (LLI) N.o. Tiles, N.o. Slices
Reset Reset 0, 0
X, Y, Z XGate, YGate, ZGate 0, 0
H HGate, RotateSingleCellPatch 2, 3
CNOT BellBasedCNOT Route-dependent, 2
T/$\mathrm{T}^{\dagger}$ MagicStateRequest, BellBasedCNOT, SinglePatchMeasurement Route-dependent, 2
S/$\mathrm{S}^{\dagger}$ YStateRequest, BellBasedCNOT, HGate, RotateSingleCellPatch Route-dependent, 10
@@ H HGate, RotateSingleCellPatch 2, 3
The first stage of lattice surgery compilation maps the Clifford+T gate set into an abstract instruction set which acts as an intermediate representation within the fault-tolerant layer. This instruction set is not to be confused with the local lattice surgery instruction set that will compose the final compiled output; the key difference between them is that the intermediate representation acts on logical qubit (patch) IDs while the local instructions act on tiles. Being an extension of what were called logical lattice instructions (LLI) in Ref. (Watkins et al., 2024), we will continue use the term LLI to refer to this intermediate representation for the sake of continuity. Due to the nature of lattice surgery compilation requiring the preparation of ancillae and usage of resource states that were not present in the original logical circuit, new patch IDs need to be generated throughout this stage of compilation. These patch IDs remain virtual until bound with patches instantiated on tiles in the second stage of compilation. It is worth noting here that only a logical dependency graph of LLI results from the first stage of compilation, while the product of the second stage is LLI organized onto logical time-slices (sliced LLI) with their compilation into local instructions given in brackets (if applicable). The former depends on dependencies of LLI alone while the latter depends upon the availability of routes, resource states, and the busy status of patches.
@@
We will now overview how each of the logical gates mentioned in the previous section map into LLI (or sequences of LLI) in the first stage of compilation (summarized in Table 1). Reset and Pauli X, Y, and Z are themselves treated as LLI. CNOT gates are simply compiled into BellBasedCNOT LLI, which act on the control and target patch IDs. We note that, while local compilation of the BellBasedCNOT LLI requires the preparation and subsequent measurement of many new logical qubit patches, we do not assign patch IDs to them because they do not have any impact on the logical dependency graph and thus their assignment is unnecessary. H gates are themselves treated as LLI but are followed by a RotateSingleCellPatch LLI. S/$\mathrm{S}^{\dagger}$ gates begin with a YStateRequest LLI acting on a newly generated patch ID (later to be associated with the nearest patch containing a Y state). Then, the appropriate sequence of LLI to execute the required H and CNOT gates according to the catalytic S gate circuit in Ref. (Fowler et al., 2012) are added. Lastly, T/$\mathrm{T}^{\dagger}$ gates begin with a MagicStateRequest LLI acting on a newly generated patch ID (later to be associated with the nearest patch containing a magic state) followed by the BellBasedCNOT LLI and SinglePatchMeasurement LLI required to implement the gate teleportation circuit from Ref. (Fowler et al., 2012). The LLI required by the corrective S gate then follow.
@@ 2.2.3. Further Processing of Intermediate Representation
The second stage of lattice surgery compilation involves the routing and slicing of the LLI resulting from the first stage. In order to translate the position-agnostic intermediate representation into instructions acting on hardware tiles, the compiler must allocate an appropriate space-time volume to each instruction while avoiding resource conflicts. The layout, which is configurable by input to the compiler (Fig. 1), provides a set of constraints at the outset of compilation. Subsequently processed LLI modify these constraints by instantiating patches on tiles along routes, binding or un-binding resource states with IDs, and/or changing the active status of patches and their edges during lattice surgery operations. Once no further LLI can be compiled onto a slice the compiler advances to the next slice, which can also modify the constraints by e.g. re-setting tiles after measurements or re-supplying magic states according to the compilation settings. Given our input instruction set and intermediate representation, our compiler binds patch IDs to available magic states and Y states, allocates busy tiles for RotateSingleCellPatch LLI, and compiles BellBasedCNOT LLI into local lattice surgery instructions in the second compilation stage.
Figure 2. The path corresponding to a BellBasedCNOT LLI
(see Fig. 4 for layout description).
Figure 3. The same as Fig. 3 except with a second BellBasedCNOT LLI whose route avoids the first.
We use results from Ref. (Beverland et al., 2022a) to compile BellBasedCNOT LLI into local instructions (see Table 2). In that protocol, the preparation and usage of a long-range Bell state requires the usage of every tile located along a free route between the control and target patches for two logical time-slices (see Fig. 3 for an example of a BellBasedCNOT route), which involves two layers of lattice surgery operations to produce local Bell states and perform Bell measurements between their end-points [^12] The specific local instructions that the BellBasedCNOT are compiled to depend on whether there is an even or odd number of tiles along the route between the control and target patches, according to Fig. 19a and c in Ref. (Beverland et al., 2022a).. While this method involves more local lattice surgery operations than do CNOT compilation protocols involving extended ancilla patches (Horsman et al., 2012; Litinski and von Oppen, 2018), it does not have increased resource overhead relative to them. Implementation details for this protocol will be given in Sec. 2.3. We re-iterate that our intention in compiling to a local instruction set rather than using extended ancilla patches is that it facilitates a simple reduction into primitives that allows for easier end-to-end compilation and circuit verification (LeBlond et al., 2023a).
@@ 2.2.4. Contention for Space
Operations that are logically parallel may not be parallelizable within the fault-tolerant layer due to geometrical constraints. For example, a CNOT gate between two qubits requires an ancilla path between the control and target, a T gate requires an ancilla path from a logical qubit to a magic state, and a Hadamard gate requires an uninitialized neighboring ancilla tile so that a patch rotation may be performed. See Fig. 3 for a simple example of resource contention between two CNOT gates. Ref. (Beverland et al., 2022a) has proposed one method for retaining logical parallelism in the fault-tolerant layer by finding maximal edge disjoint path sets. Among other things, they found that a layer of $k$ parallel CNOT gates needs depth $\Omega(\sqrt{k})$ to execute on a surface code architecture in the worst-case. In our compilation strategy, we exploit logical parallelism within the fault-tolerant layer by relying on a simple heuristic algorithm (A*) to find vertex disjoint paths and greedily allocating these paths to slices (see Algorithm 1 in Sec. 2.3 for details). In Sec. 3 we will show that, in our implementation, the number of slices scales similarly to these expectations from Ref. (Beverland et al., 2022a).
@@ 2.2.5. Choice of Surface Code Layout
Within a lattice surgery paradigm, surface code patches can be simplistically thought of as rectangular entities that are merged and split through edge-wise interactions to generate entanglement or transport quantum information. This results in a type of quantum circuit layout problem, where in general the layout of qubits on a constrained hardware architecture is optimized to reduce circuit depth (Paler et al., 2023). In surface code compilation, layouts of equally-sized tiles on a two-dimensional grid are often assumed (Litinski, 2019a; Beverland et al., 2022a), though it is not always beneficial to have only one tile size available throughout a computation (Litinski, 2019b; Beverland et al., 2022b). Equal X and Z code distances are also not necessarily beneficial in the realistic setting of biased noise (Chamberland and Campbell, 2022b). In proposals for quantum computer architectures, there is a trade-off between compression of layouts in space and the accessibility of surface code patch edges, as compression in space can lead to overhead in time due to the need to rotate patches to expose needed operators (Litinski, 2019a). While keeping the number of tiles dedicated to routing space minimal under the condition that all patch edges are exposed is generally considered to be desirable in sequential Pauli-based computation (Litinski, 2019a; Beverland et al., 2022b; Chamberland and Campbell, 2022b), it is not as clear in the case of direct Clifford+T compilation where liberal routing space can aid in parallel compilation of logically parallel gates (Beverland et al., 2022a). In this paper, we take the definition of logical tile presented in Ref. (LeBlond et al., 2023a) and make use of a family of layouts related to the one from Ref. (Beverland et al., 2022a) [^13] We will call this family of layouts EDPC layouts because of the paper that proposed them.. We assume that tile sizes are uniform within lattice surgery compilation since, while we do use two levels of distillation with two different code distances, we do not explicitly compile distillation circuits (as explained in Sec. 4).
@@ 2.3. Our Revisions to the Lattice Surgery Compiler
Much of the functionality needed to implement our compilation strategy from Sec. 2.2 did not originally exist in LSC. However, we leveraged the extensibility of the LSC framework to improve layout generation including the handling of magic states, expand the options for processing logical gates into the intermediate representation (LLI), create a compilation pipeline to handle DAG-based parallelism in the LLI layer, and implement a new compilation layer for processing LLI into local instructions. In this section, we will provide details about our revisions to the LSC in a few areas where we feel that more specificity is warranted than what was presented in Sec. 2.2. We will first review some of the functionality of the original LSC from Ref. (Watkins et al., 2024) in order to provide context for our contributions to it.
@@ 2.3.1. Lattice Surgery Compiler Review
  • Compilation is implemented as a streamed process, so that compilation could be performed in real-time based on input to classical control software while the quantum computer executes previously compiled instructions. The streaming pipeline (which we retained as an option in our revised version) has the benefit of being able to compile circuits that are too large to fit in memory. It also allows for a limited form of parallelism by compiling streamed instructions onto the same slice until it encounters one that cannot be routed. However, the parallelism is limited by the lack of a logical dependency graph prior to time-slicing instructions, and by multi-stage instructions which automatically request new time-slices for follow-up instructions.
@@ 2.3.2. Layout Generation
We have improved the layout generation capabilities within the LSC by implementing three new types of tiles (specified by text characters as described previously). In the new version of the compiler, there are specified tiles that are reserved for magic states (labeled ‘M’) on which magic states are replenished according to the disttime and nostagger parameters, which respectively specify the rate at which magic states are replenished at these tiles and whether or not they are replenished in a staggered fashion. This improvement enables compilation of T/$\mathrm{T}^{\dagger}$ gates to be dissociated from the space-time footprint of magic state distillation, allowing much greater flexibility in resource estimation and optimization. The post-hoc resource analysis discussed in Sec. 4 accounts for both the space-time footprint of magic state factories and the cost to maintain distilled state storage banks, both of which are determined after the cadence of sliced MagicStateRequest LLI is known from the LSC output. Although magic state distillation is not specifically compiled in our framework, our analysis of magic state management (generation, storage, and consumption) makes no approximations as to the availability and consumption of these states at any point in the computation. The second new type of tile we have implemented are tiles that store pre-distilled Y states (labeled ‘Y’). (Previously, they were produced on-the-fly assuming a twist-based protocol (Watkins et al., 2024), but we have chosen to employ a twist-free instruction set here). Finally, dead tiles (tiles on which no operations can be performed and which are avoided by the router) have also been implemented (labeled ‘X’) so that non-rectangular layouts can be specified (though we have not yet made use of this).
@@
For our resource estimates, we implemented a layout generator that constructs the EDPC layout as illustrated in Fig. 4, accessible through the -L edpc command line option (Fig. 4(a) represents the EDPC layout that is generated by default using this setting). In Fig. 4, the brown tiles in the bulk (bound to patch IDs) correspond to patches encoding logical qubits from the input circuit. Blue and pink/purple tiles on the boundary correspond to pre-distilled Y states and tiles that are reserved for magic states, respectively. While these have no patch IDs by default, they are bound with IDs upon encountering YStateRequest and MagicStateRequest LLI during the second stage of compilation. White tiles indicate ancilla tiles that can be used for routing (in the case of our compilation strategy, this is where long-range Bell states are prepared and where patch rotations take place). In order to observe the effects of different data:ancilla tile ratios, we implemented num_lanes and condensed options to -L edpc. The effects of these options on the generated layout are depicted in Fig. 4(b) and (c), respectively. The former option can be used to open up more lanes for routing CNOT gates at the expense of greater spatial overhead, while the latter option decreases the total routing space by blocking off two edges of each logical qubit. The effects of these options on circuit depth and space-time volume will be explored in Sec. 3.
@@
Figure 4. EDPC layouts for 9 logical qubits, generated using the (a) default, (b) num_lanes = 2, and (c) condensed settings. Brown tiles in the bulk correspond to logical qubits. Blue and pink/purple tiles on the boundary correspond to Y states and tiles reserved for magic states, respectively. White tiles indicate tiles dedicated to routing instructions and the preparation of ancillae.